OpenStudy (lemanual):

A normally-distributed data set has a mean of 30 and a standard deviation of 5. What is the area under the standard normal curve for data greater than 35? Enter your answer in the box. Round your answer as a decimal rounded to the nearest thousandth. Note: Do not give your answer as a percent.

OpenStudy (welshfella):

Hold on - I'm stretching my powers of memory now!!!

OpenStudy (lemanual):

ok lol

OpenStudy (welshfella):

work out the z score first that is (35 - mean ) / std dev = (35 - 30 ) / 5

OpenStudy (lemanual):

I'm confused :/

OpenStudy (lemanual):

Sorry :/

OpenStudy (welshfella):

35-30 = 5 5/5 = 1 z score = 1

OpenStudy (welshfella):



OpenStudy (welshfella):

the area with z> 1 is the shaded area

OpenStudy (welshfella):

this is equals the area corresponding to data greater than 35

OpenStudy (welshfella):

You can find this area from a table of a normal distribution

OpenStudy (welshfella):